In a Young's Slit experiment a wavelength of 589 nm is used and intensity maxima and minima are observed on screen one metre from the slits. A total of 7 maxima are observed on the screen, with the outer lying at +-1.886 m from the central axis. What is the slit separation?
Answer #1

here the conditin is y = m*lamda*L/d

y = distance form central maximum to 7th central maximum

lamda = wavelength of light used = 589 nm

m = 7

L = 1 m

d = slit seperation

d = m*lamda*L/y = (7*589*10^-9*1)/(1.886) = 2.186*10^-6 m = 2.186 micro meter